# Non-degenerate mixed maps and contact structures

### v1

@@ 3.2. Links of mixed ICIS
    In [28, Theorem 1], the mixed function is the pullback of non-degenerate convenient holomorphic functions. As we have seen, these two conditions together are generalized by partial non-degeneracy, which is also sufficient to obtain the same result in the case $k=1$.
@@
    The first paragraph in the previous proof implies that a mixed link determined by a mixed ICIS is a positive (respectively, negative) contact submanifold of the sphere $\mathbb{S}_{r}^{2n-1}\subset\mathbb{C}^{n}$ if and only if the associated mixed map is (resepctively, anti-)holomorphic-like.
  3. (3)
    We shall refer to the contact structure above as natural and denote it by $\xi_{r}$.
@@
Let us denote $D_{H}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=D\circ\varphi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ the associated real function of $H$ and suppose the statement is false. Moreover, let us introduce coordinates $\varphi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$. By the Curve Selection Lemma, there exists an open neighborhood $U$ of the origin and a real analytic curve $\lambda:(0,1]\longrightarrow U\subset\mathbb{C}^{n}$ such that:
@@
$\begin{cases}&\lim_{t\to 0}\lambda(t)=0\Longrightarrow\lim_{t\to 0}\varphi(\lambda(t))=0,\\ &H(z(t),{}\mkern 3.0mu\overline{\mkern-3.0muz}(t))\equiv 0\Longrightarrow G(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0,\\ &D_{H}(z(t),{}\mkern 3.0mu\overline{\mkern-3.0muz}(t))\equiv 0\Longrightarrow D(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0,\end{cases}$
@@
where $z(t),w(t)$ denote the restrictions of the coordinates to the curve $\lambda(t)$. Then $\varphi(\lambda(t))$ is a real analytic curve on $V_{G}$ on which the link $K_{G}=V_{G}\cap\mathbb{S}_{r}^{2n-1}$ of $g$ is not a contact submanifold for all $r>0$ such that $\mathbb{S}_{r}^{2n-1}\subset U$. But this is a contradiction with the initial assumption. ∎
Remark 3.12.
Notice that the existence of the natural contact structure on a link defined by a (anti-)holomorphic-like mixed map only makes sense if the ambient space has a holomorphic structure.
@@
This proposition allows us the following conclusion. Let $(V,x)\subset(M,m)$ be the germ of a mixed variety with an isolated singularity at $x$ and $(M,m)$ the germ of a complex analytic manifold. If there exists a coordinate system $z$ of $M$ for which $K_{V}=V\cap\mathbb{S}_{r,z}^{2n-1}$ is a contact manifold, where $\mathbb{S}_{r,z}^{2n-1}$ is the usual sphere on the coordinate system $z$, then the same assertion holds for any other holomorphic local coordinate system of $M$.
@@ 3.3. Open books
We prove in this section the existence of open books adapted to the natural contact structures on mixed links of ICIS. Given a mixed function $g$ with an isolated singularity at the origin, we establish an extra condition related to the mixed ICIS $G=\phi^{*}F$ that allows us to derive an open book that is further adapted. This hypothesis is based on the proof of [5, Proposition 3.2], in the holomorphic non-isolated singularity context, and [28, Theorem 4].
@@
By [28, p.266], the expression of the Reeb vector field $R_{c}$ of $\alpha_{c}$ becomes:
@@
Suppose that $g=\phi^{*}f$, where $\phi$ is a homogeneous mixed covering and $f$ a holomorphic function. By [28, Lemma 4], $\|v_{1}\|^{2}\geq\|v_{2}\|^{2}$ and the equality holds if and only if $\nabla\Theta_{g}=\lambda R$ for some $\lambda\in\mathbb{C}$, where
@@
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ and $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ be partially non-degenerate holomorphic map and function germs, respectively. Let $\phi_{a,b}$ and $\phi_{c,d}$ be homogeneous mixed coverings, where $a>b$ and $c>d$, and define the pullbacks $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{a,b}^{*}F$ and $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{c,d}^{*}f$. Suppose further that $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ defines with $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ a mixed ICIS germ $\Psi:=(G,g):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k+1},0)$. Then the restriction
@@
First, since the map $\Psi=(G,g):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k+1},0)$ is an isolated complete intersection singularity, there exist $r_{0}>0$ and $\eta>0$ such that $\Psi^{-1}(s,t)$ intersect the sphere $\mathbb{S}^{2n-1}_{r}$ transversely for all $r<r_{0}$ and $\|(s,t)\|<\eta$. Whence, the fibers $g^{-1}(t)$ intersect $K_{G}$ transversely for $t$ sufficiently small. Recall that $g$ has an isolated singularity at the origin by Proposition 2.3. By Lemma 3.3, this implies that $\Theta_{g}$ defines an open book in $K_{G}$. On the other hand, by Theorem 3.9, the link $K_{g}=V_{g}\cap\mathbb{S}^{2n-1}_{r}$ is a contact submanifold as well as its restriction to $K_{G}$. Considering our convention for the orientations, it remains to verify that the fibers of $\Theta_{g}$ have the natural symplectic structure. We shall apply the same strategy of [6, Theorem 3.9] and [28, Theorem 4]. That is, we consider the modification $\alpha_{c}$ in (3.4) which induces the same hyperplane distribution but satisfies $d\Theta(R_{c})>0$. Define
@@
For the first case, it is enough to choose a sufficiently large $c>0$ to make $d\Theta_{g}(R_{c})>0$. Moreover, we have claimed that $\|v_{1}\|=\|v_{2}\|$ if and only if $\nabla\Theta_{g}(w)=\lambda R(w)$ and, in this case, $d\Theta_{g}(R)=\operatorname{Re}\lambda\|R\|^{2}$. By [28, Lemma 5], if $w\in Z_{\delta}$ is a solution for this equation, then $\operatorname{Re}\lambda>0$. We conclude that $d\Theta_{g}(R)>0$. ∎
@@ 4. Natural and Milnor fillable structures
Let $G:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed ICIS germ and $(V,0)\subset\mathbb{C}^{n}$ be a complex germ with an isolated singularity at the origin. Let $K_{G}$ be the mixed link and suppose the existence of a map germ $\varphi:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ which is a diffeomorphism on $K_{V}$. One can define a contact structure on $K_{G}$ by setting $\xi_{G}=d\phi(\xi_{V})$. This occurs for mixed Hamm ICIS in Theorem 2.12. See also [26] and [15].
If there exists another diffeomorphism $\psi:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ and we set a contact structure $\xi_{G}^{\prime}=d\psi(\xi_{V})$ on $K_{G}$ induced by $\psi$, it is clear that $\xi_{G}^{\prime}$ and $\xi_{G}$ are contactomorphic. Moreover, if $(\Theta,N)$ is an open book adapted to $\xi_{V}$, the induced one is $\psi^{*}\Theta$, where $\psi=\varphi^{-1}$ and the binding is $\phi(N)$. Recall that we have set $\alpha$ as the contact form (3.1) on the sphere.
@@
since $\mkern 3.0mu\overline{\mkern-3.0mu\partial}h=\partial{}\mkern 3.0mu\overline{\mkern-3.0muh}=0$ because $h$ is holomorphic. We obtain that
@@
$d(\Theta\circ\psi)_{z}(R)=\sum_{i,j}$ $\frac{\partial h}{\partial z_{j}}\frac{z_{j}}{h}\left[\frac{\partial}{\partial z_{j}}\left(\frac{\psi_{i}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2}\right)-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\left(\frac{\psi_{i}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2}\right)\right]$
- $\frac{\partial{}\mkern 3.0mu\overline{\mkern-3.0muh}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\frac{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}{{}\mkern 3.0mu\overline{\mkern-3.0muh}}\left[\frac{\partial}{\partial z_{j}}\left(\frac{\psi_{i}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2i}\right)-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\left(\frac{\psi_{i}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{i}}{2i}\right)\right].$
If $\alpha_{j}=\frac{\partial h}{\partial z_{j}}\frac{z_{j}}{h}$, we may rewrite the expression as

### v3

@@ 3.2. Links of mixed ICIS
    The formulation in the previous proof implies that a mixed link determined by a mixed ICIS is a positive (respectively, negative) contact submanifold of the sphere $\mathbb{S}_{r}^{2n-1}\subset\mathbb{C}^{n}$ if and only if the associated mixed map is (respectively, anti-)holomorphic-like. We shall refer to the contact structure above as natural and denote it by $\xi_{r}$.
@@
    Since this structure does not depend on the radius of the sphere, the same holds for the mixed links. Moreover, for any other mixed map germ $G=(g^{1},\dots,g^{n}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ that defines the same mixed ICIS $V_{F}$, the contact condition $\alpha\wedge(d\alpha)^{n-(k+1)}\neq 0$ can be replaced by (3.3) in terms of $\mathcal{D}^{G}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ and the result does not depend on the defining equations.
@@
Let us denote $\mathcal{D}^{H}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\mathcal{D}^{G}\circ\varphi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ the associated analytic function of $H$ and suppose the statement is false. Moreover, let us introduce coordinates $\varphi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$. By the Curve Selection Lemma, there exists an open neighborhood $\mathring{\mathbb{B}}_{r}$ of the origin and a real analytic curve $\lambda:(0,1]\longrightarrow\mathring{\mathbb{B}}_{r}\subset\mathbb{C}^{n}$ such that:
@@
$\begin{cases}&\lim_{t\to 0}\lambda(t)=0\Longrightarrow\lim_{t\to 0}\varphi(\lambda(t))=0,\\ &H(z(t),{}\mkern 3.0mu\overline{\mkern-3.0muz}(t))\equiv 0\Longrightarrow G(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0,\\ &\mathcal{D}^{H}(z(t),{}\mkern 3.0mu\overline{\mkern-3.0muz}(t))\equiv 0\Longrightarrow\mathcal{D}^{G}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0,\end{cases}$
@@
where $z(t),w(t)$ denote the restrictions of the coordinates to the curve $\lambda(t)$. Then $\varphi(\lambda(t))$ is a real analytic curve on $V_{G}$ on which the link $K_{G}=V_{G}\cap\mathbb{S}_{r^{\prime}}^{2n-1}$ of $g$ is not a contact submanifold for all $r^{\prime}>0$ such that $\mathbb{S}_{r^{\prime}}^{2n-1}\subset\mathring{\mathbb{B}}_{r}$. But this is a contradiction with the initial assumption. ∎
@@
Notice that the existence of the natural contact structure on a link defined by a (anti-)holomorphic-like mixed map only makes sense if the ambient space has a holomorphic structure. This proposition allows us to draw the following conclusion. Let $(V,x)\subset(M,m)$ be the germ of a mixed variety with an isolated singularity at $x$ and $(M,m)$ the germ of a complex analytic manifold. If there exists a coordinate system $z$ of $M$ for which $K_{V}=V\cap\mathbb{S}_{r,z}^{2n-1}$ is a contact manifold, where $\mathbb{S}_{r,z}^{2n-1}$ is the usual sphere on the coordinate system $z$, then the same assertion holds for any other holomorphic local coordinate system of $M$. In the sequel, we investigate conditions under which natural contact structures exist on mixed links.
Definition 3.12.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ. We say that $F$ is contact non-degenerate if for any vector $P$ of positive integers and $p\in\mathbb{C}^{*n}$ such that $F_{P}(p)=0$, we have that $\mathcal{D}^{F_{P}}(p,{}\mkern 3.0mu\overline{\mkern-3.0mup})\neq 0$.
Proposition 3.13.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a convenient mixed map germ. Then the following statements hold:
  1. (1)
    If $F$ is contact degenerate, then the restriction $F^{I}$ is also contact non-degenerate for every $I\subset\{1,\dots,n\}$ such that $|I|>k$ and $V_{F^{I}}$ has positive dimension.
  2. (2)
    If $F$ is contact non-degenerate, then it is strictly (anti-)holomorphic-like.
  3. (3)
    If $F$ is holomorphic, non-degenerate, and strongly convenient, then it is contact non-degenerate.
Proof.
The first item of Proposition 2.9 implies that under the convenience condition, the face functions $F^{I}_{P}$ of the restriction of a map $F$ can be identified with face functions of $F$ itself. Therefore, we obtain the first assertion. If the second statement is false, by the Curve Selection Lemma, there exists an analytic curve $w:(0,1]\longrightarrow\mathbb{C}^{n}$ such that $\lim_{t\to 0}w(t)=0$, $w(t)\subset V_{F}$, and $\mathcal{D}^{F}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0$. Let $I=\{i:w_{i}(t)\not\equiv 0\}$. By Equation (2.1), we know that the partial derivatives of the coordinate functions $f^{l,I}$ evaluated at $(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$ have analytic expansions whose the lower degree terms are the derivative of the face functions $f^{l,I}_{P}$, where $P$ is a vector of positive integers determined by the analytic expansion of $w(t)$. For each index $l$ and pair $i,j\in I$, we have that:
$\mathcal{B}_{i,j}^{f^{l,I}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))=\mathcal{B}_{i,j}^{f^{l,I}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{2d_{l}-p_{i}-p_{j}}+o(t).$
Thus, for each set $J=\{j_{1},\dots,j_{k+1}\}$ of distinct indices and order $k+1$, it follows that:
$\mathcal{D}^{F^{I}}_{J}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))=\mathcal{D}^{F_{P}^{I}}_{J}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{D_{J}}+o(t),$
where $D_{J}=p_{j_{1}}+p_{j_{2}}+2d_{1}-p_{j_{3}}-p_{j_{4}}+\dots+2d_{k}-p_{j_{2k+1}}-p_{j_{2k+2}}$. Let $D_{1},\dots,D_{m}$ be the possible total distinct degrees among all sets $J$ of order $k+1$. Observe that
$\mathcal{D}^{F^{I}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))=\left(\sum_{D_{J}=D_{1}}\mathcal{D}_{J}^{F_{P}^{I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})\right)t^{D_{1}}+\dots+\left(\sum_{D_{J}=D_{m}}\mathcal{D}_{J}^{F_{P}^{I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})\right)t^{D_{m}}+o(t)\equiv 0,$
where $o(t)$ denotes terms of higher order than $\max_{i}\{D_{i}\}$ in $t$. We obtain that:
$\mathcal{D}^{F_{P}^{I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=\sum_{J}\mathcal{D}_{J}^{F_{P}^{I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=\left(\sum_{D_{J}=D_{1}}\mathcal{D}_{J}^{F_{P}^{I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})\right)+\dots+\left(\sum_{D_{J}=D_{m}}\mathcal{D}_{J}^{F_{P}^{I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})\right)=0.$
This contradicts the contact non-degeneracy of $F^{I}$ in the first item. In the last case, for every vector $P$ of positive integers, the map $F_{P}$ is a non-degenerate and convenient holomorphic map germ. By Theorem 2.13, this defines a holomorphic ICIS. Since these varieties define links endowed with the natural contact structure, we conclude that $\mathcal{D}^{F_{P}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\neq 0$ in a small neighborhood of the origin. ∎
The next result shows that mixed maps germs that are higher order deformations of holomorphic ICIS define links that admit the natural contact structure.
Theorem 3.14.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ of the following form:
$F(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=G(z)+H(z,{}\mkern 3.0mu\overline{\mkern-3.0muz}),$
where $G:(g^{1},\dots,g^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is a non-degenerate and strongly convenient holomorphic map germ such that $\Gamma(f^{j})=\Gamma(g^{j})$ for every $j=1,\dots,k$, and $H(z,{}\mkern 3.0mu\overline{\mkern-3.0muz}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is a mixed map germ. Then $F$ is contact non-degenerate. In particular, the link $K_{F}=V_{F}\cap\mathbb{S}_{r}^{2n-1}$ is a contact submanifold for every sufficiently small $r>0$.
Proof.
We observe first that $F(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is a mixed ICIS, since its Newton polyhedron is that of $G$, which is a holomorphic ICIS. Secondly, notice that for each vector $P$ of positive integers and $I\subset\{1,\dots,n\}$, the functions $\mathcal{D}^{F^{I}_{P}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ and $\mathcal{D}^{G^{I}_{P}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ coincide, provided that $g^{j,I}_{P}=f^{j,I}_{P}$ for every $j$. This implies that $F$ is contact non-degenerate by the third item of Proposition 3.13. The conclusion is a consequence of Theorem 3.9. ∎
@@ 3.3. Open books
We prove in this section the existence of open books adapted to the natural contact structures on mixed links of ICIS. Given a mixed function $g$ with an isolated singularity at the origin, we establish an extra condition related to the mixed ICIS $G=\phi^{*}F$ that allows us to derive an open book that is further adapted. This hypothesis is based on the proof of [7, Proposition 3.2], in the holomorphic non-isolated singularity context, and [32, Theorem 4].
@@
By [32, p.266], the expression of the Reeb vector field $R_{c}$ of $\alpha_{c}$ becomes:
@@
Suppose that $g=\phi^{*}f$, where $\phi$ is a homogeneous mixed covering and $f$ a holomorphic function. By [32, Lemma 4], $\|v_{1}\|^{2}\geq\|v_{2}\|^{2}$ and the equality holds if and only if $\nabla\Theta_{g}=\lambda R$ for some $\lambda\in\mathbb{C}$, where
@@
Let $G:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a strictly holomorphic-like mixed map germ as in Theorem 3.9 and $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ a non-degenerate convenient holomorphic function germ. Let $\phi_{c,d}$ be a homogeneous mixed covering, where $c>d$, and define the pullback $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{c,d}^{*}f$. Suppose further that $g(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ defines with $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})$ a mixed ICIS germ $\Psi:=(G,g):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k+1},0)$. Then the restriction
@@
First, since the map $\Psi=(G,g):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k+1},0)$ is an isolated complete intersection singularity, there exist $r_{0}>0$ and $\eta>0$ such that $\Psi^{-1}(s,t)$ intersect the sphere $\mathbb{S}^{2n-1}_{r}$ transversely for all $r<r_{0}$ and $\|(s,t)\|<\eta$. Whence, the fibers $g^{-1}(t)$ intersect $K_{G}$ transversely for $t$ sufficiently small. Recall that $g$ has an isolated singularity at the origin by Proposition 2.3. By Lemma 3.3, this implies that $\Theta_{g}$ defines an open book in $K_{G}$. On the other hand, by Theorem 3.9, the link $K_{g}=V_{g}\cap\mathbb{S}^{2n-1}_{r}$ is a contact submanifold as well as its restriction to $K_{G}$. Considering our convention for the orientations, it remains to verify that the fibers of $\Theta_{g}$ have the natural symplectic structure. We shall apply the same strategy of [6, Theorem 3.9] and [32, Theorem 4]. That is, we consider the modification $\alpha_{c}$ in (3.4) which induces the same hyperplane distribution but satisfies $d\Theta(R_{c})>0$. Define
@@
For the first case, it is enough to choose a sufficiently large $c>0$ to make $d\Theta_{g}(R_{c})>0$. Moreover, we have claimed that $\|v_{1}\|=\|v_{2}\|$ if and only if $\nabla\Theta_{g}(w)=\lambda R(w)$ and, in this case, $d\Theta_{g}(R)=\operatorname{Re}\lambda\|R\|^{2}$. By [32, Lemma 5], if $w\in Z_{\delta}$ is a solution for this equation, then $\operatorname{Re}\lambda>0$. We conclude that $d\Theta_{g}(R)>0$. ∎
@@ 4. Natural and Milnor fillable structures
Let $G:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed ICIS germ and $(V,0)\subset\mathbb{C}^{n}$ be a complex germ with an isolated singularity at the origin. Let $K_{G}$ be the mixed link and suppose the existence of a map germ $\phi:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ which is a diffeomorphism on $K_{V}$. One can define a contact structure on $K_{G}$ by setting $\xi_{G}=d\phi(\xi_{V})$. This occurs for mixed Hamm ICIS in Theorem 2.24. See also [30] and [17].
If there exists another map $\phi^{\prime}:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ whose restriction to $K_{V}$ defines a diffeomorphism with $K_{G}$ and we set a contact structure $\xi_{G}^{\prime}=d\phi^{\prime}(\xi_{V})$ on $K_{G}$ induced by $\phi^{\prime}$, it is clear that $\xi_{G}^{\prime}$ and $\xi_{G}$ are contactomorphic. Moreover, if $(\Theta,N)$ is an open book adapted to $\xi_{V}$, the induced one is $\psi^{*}\Theta$, where $\psi=\phi^{-1}$ and the binding is $\phi(N)$. Recall that we have set $\alpha$ as the contact form (3.1) on the sphere. In what follows, we may relax this construction by supposing that the smooth map $\phi:(\mathbb{C}^{n},K_{V})\longrightarrow(\mathbb{C}^{n},K_{G})$ restricts to the links as a diffeomorphism.
@@
since $\mkern 3.0mu\overline{\mkern-3.0mu\partial}h=\partial{}\mkern 3.0mu\overline{\mkern-3.0muh}=0$ because $h$ is holomorphic. Putting $\psi=(\psi_{1},\dots,\psi_{n})$ and writing the matrix $d\psi$ in the coordinates $(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$, we obtain that:
$(d\psi\circ R)_{z}=\begin{pmatrix}\frac{\partial\operatorname{Re}\psi_{1}}{\partial z_{1}}&\frac{\partial\operatorname{Re}\psi_{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}&\dots&\frac{\partial\operatorname{Re}\psi_{1}}{\partial z_{n}}&\frac{\partial\operatorname{Re}\psi_{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}}\\ \frac{\partial\operatorname{Im}\psi_{1}}{\partial z_{1}}&\frac{\partial\operatorname{Im}\psi_{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}&\dots&\frac{\partial\operatorname{Im}\psi_{1}}{\partial z_{n}}&\frac{\partial\operatorname{Im}\psi_{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}}\\ \vdots&\vdots&\vdots&\cdots&\vdots\\ \frac{\partial\operatorname{Re}\psi_{n}}{\partial z_{1}}&\frac{\partial\operatorname{Re}\psi_{n}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}&\dots&\frac{\partial\operatorname{Re}\psi_{n}}{\partial z_{n}}&\frac{\partial\operatorname{Re}\psi_{n}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}}\\ \frac{\partial\operatorname{Im}\psi_{n}}{\partial z_{1}}&\frac{\partial\operatorname{Im}\psi_{n}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}&\dots&\frac{\partial\operatorname{Im}\psi_{n}}{\partial z_{n}}&\frac{\partial\operatorname{Im}\psi_{n}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}}\end{pmatrix}\cdot\begin{pmatrix}z_{1}\\ -{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\\ \vdots\\ z_{n}\\ -{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n}\end{pmatrix}=\begin{pmatrix}\sum_{i}\frac{\partial\operatorname{Re}\psi_{1}}{\partial z_{i}}z_{i}-\frac{\partial\operatorname{Re}\psi_{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\\ \sum_{i}\frac{\partial\operatorname{Im}\psi_{1}}{\partial z_{i}}z_{i}-\frac{\partial\operatorname{Im}\psi_{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\\ \vdots\\ \sum_{i}\frac{\partial\operatorname{Re}\psi_{n}}{\partial z_{i}}z_{i}-\frac{\partial\operatorname{Re}\psi_{n}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\\ \sum_{i}\frac{\partial\operatorname{Im}\psi_{n}}{\partial z_{i}}z_{i}-\frac{\partial\operatorname{Im}\psi_{n}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\end{pmatrix}$
Hence, the expression becomes:
@@
$d(\Theta\circ\psi)_{z}(R)=\sum_{i,j}$ $\frac{\partial h}{\partial z_{j}}\frac{1}{h}\left[\frac{\partial}{\partial z_{i}}\left(\frac{\psi_{j}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{j}}{2}\right)z_{i}-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\left(\frac{\psi_{j}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{j}}{2}\right){}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\right]$
- $\frac{\partial{}\mkern 3.0mu\overline{\mkern-3.0muh}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\frac{1}{{}\mkern 3.0mu\overline{\mkern-3.0muh}}\left[\frac{\partial}{\partial z_{i}}\left(\frac{\psi_{j}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{j}}{2{\bf i}}\right)z_{i}-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\left(\frac{\psi_{j}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{j}}{2{\bf i}}\right){}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\right].$
If we set $h=z_{1}$, then: